OpenStudy (anonymous):

sin4theta = cos2theta please solve for -pi ≤ theta ≤ pi

OpenStudy (campbell_st):

is it \[\sin(4\theta) = \cos(2 \theta)\] or \[\sin^4(\theta) = \cos^2{\theta}\]

OpenStudy (anonymous):

the first one: sin(4θ)=cos(2θ)

OpenStudy (campbell_st):

ok use the double angle expansion \[\sin(2(2\theta)) = 2\sin(2\theta)\cos(2\theta)\] than then you have \[2\sin(2\theta)\cos(2\theta) = \cos(2\theta)\] this should make it easier...

OpenStudy (campbell_st):

divide both sides by \[\cos(2 \theta)\]

OpenStudy (anonymous):

\[\sin 4\theta = \sin(2\theta+2\theta) = 2\sin(2\theta)\cos(2\theta)\ = \cos(2\theta)\] cancel the \[\cos2\theta\] from both sides and you got \[\sin2\theta = \frac{ 1 }{ 2 } 2\theta = 30 and \theta = 15°\]

OpenStudy (anonymous):

thx so much for all your help :D

OpenStudy (campbell_st):

the only problem is that's not the complete solution, you need to check the defined domain.

OpenStudy (anonymous):

?? defined domain? btw doesn't dividing both sides by cos2θ make you lose some roots?

OpenStudy (anonymous):

@alexnhchong You are losing the possible root of θ when cos(2θ) = 0

OpenStudy (anonymous):

So you can always just check that if you're concerned.

OpenStudy (anonymous):

ic. again, thx a lot to those who helped :)